A ball is thrown vertically upward with a speed of 15 m/s. Determine the time of flight when it
returns to its original position.
1
Expert's answer
2021-10-03T17:05:02-0400

"u = 15 m\/s" "v = 0 m\/s"  "v = u +at" "0 = 15-10t" "t = 1.5 sec" As the ball will return to its original position, the time taken for the ball to go up will be same as that it will take to come down.  Hence total time taken = 3 sec.